OpenStudy (anonymous):

A tank is full of water. Find the work required to pump the water out of the spout. (Use 9.8 m/s2 for g. Use 1000 kg/m3 as the density of water. Assume r = 3 m and h = 1 m.)

OpenStudy (anonymous):

http://www.webassign.net/scalcet7/6-4-022-alt.gif

OpenStudy (anonymous):

any idea how to solve this?

OpenStudy (jhannybean):

 \[\large W = [(1000\cdot 9.8)\cdot \pi\cdot x^2 dy ]\cdot (4+y)\] \((1000 \cdot 9.8) \) = weight of the water x force of gravity \(x^2 \) = radius of the slab we're calculating with respect to the axis of integration. \((4+y)\) = distance the water has to move up to be pumped out of the container.